Over a distance of 120 km, the average speed of a train is 40km/h faster than that of a car. If the train covers the distance in 30min less time, find its average speed.
an equation for this pls?
thanks =D

Answer by scott8148(6628)

let x="speed of train"

from car ... t=d/r ... t=120/(x-40)

from train ... r=d/t ... x=120/((120/(x-40))-30)